OpenStudy (anonymous):

how to expand this function A(z)=1/(1-z-z^2)?????

OpenStudy (anonymous):

(1-z(1+z))^-1

OpenStudy (anonymous):

ya then what to expand this using binomial thm???

OpenStudy (anonymous):

1+z(z+1)+[z(Z+1)]^2+........

OpenStudy (anonymous):

can u specify the formula here??

OpenStudy (anonymous):

the formula is (1-x)^-1=1+x+x^2+x^3+..... here x=z(z+1)

OpenStudy (anonymous):

thanks